Given three numbers A, B and C. Construct an algorithm to compute and print out the sum, the average, and the product of these values.
1
Expert's answer
2021-09-28T01:42:12-0400
Start
  Read A, B and C
  Sum= A+B+C
  Average = Sum/3
  Product =A*B*C
  Print Sum
  Print Average 
  Print Product 
Stop